# Tree Search for Language Model Agents

### v2

@@ Abstract
Autonomous agents powered by language models (LMs) have demonstrated promise in their ability to perform decision-making tasks such as web automation. However, a key limitation remains: LMs, primarily optimized for natural language understanding and generation, struggle with multi-step reasoning, planning, and using environmental feedback when attempting to solve realistic computer tasks. Towards addressing this, we propose an inference-time search algorithm for LM agents to explicitly perform exploration and multi-step planning in interactive web environments. Our approach is a form of best-first tree search that operates within the actual environment space, and is complementary with most existing state-of-the-art agents. It is the first tree search algorithm for LM agents that shows effectiveness on realistic web tasks. On the challenging VisualWebArena benchmark, applying our search algorithm on top of a GPT-4o agent yields a 39.7% relative increase in success rate compared to the same baseline without search, setting a state-of-the-art success rate of 26.4%. On WebArena, search also yields a 28.0% relative improvement over a baseline agent, setting a competitive success rate of 19.2%. Our experiments showcase the effectiveness of search for web agents, and we demonstrate that performance scales with increased test-time compute. We conduct a thorough analysis of our results to highlight improvements from search, limitations, and promising directions for future work. Our code and models are publicly released at jykoh.com/search-agents.
@@ 1 Introduction
Building agents that can perceive, plan, and act autonomously has been a long standing goal of artificial intelligence research (Russell & Norvig, 1995; Franklin & Graesser, 1996). In recent years, the advent of large language models (LMs) with strong general capabilities has paved the way towards building language-guided agents that can automate computer tasks. However, the best LM agents today are still far worse than humans. On the realistic web benchmarks WebArena (Zhou et al., 2024b) and VisualWebArena (Koh et al., 2024), humans succeed on 78% and 89% of tasks respectively, but agents — even those powered by the latest frontier models — are far worse, typically achieving success rates below 20%. One significant bottleneck in existing agents arises from their inability to leverage test-time computation for exploration and multi-step planning. Search and planning is especially important in open ended web environments, as the potential action space (i.e., all possible actions one can take on a webpage) is much larger than in most video games or text-based simulators. There are often multiple plausible actions that must be sequenced to reach a goal, and being able to efficiently explore and prune trajectories is essential.
In artificial intelligence systems, one effective strategy for leveraging test-compute to improve results is search: iteratively constructing, exploring, and pruning a graph of intermediate states and possible solutions (Newell et al., 1959; Laird, 2019; Silver et al., 2016). The effectiveness of search algorithms has been shown time and time again, enabling models to achieve or surpass human-level performance on a variety of games, including Go (Silver et al., 2016; 2017), poker (Brown & Sandholm, 2018; 2019), and Diplomacy (Gray et al., 2020).
@@
Our experiments show that this search procedure is complementary with existing LM agents, and enables these models to perform better on harder and longer horizon tasks. On VisualWebArena (Koh et al., 2024), search improves the performance of a baseline GPT-4o (OpenAI, 2024) agent by 39.7% relative to the baseline without search, setting a new state-of-the-art (SOTA) success rate of 26.4%. On WebArena (Zhou et al., 2024b), search is also highly effective, contributing a 28.0% relative improvement (yielding a competitive success rate of 19.2%). We also demonstrate that our search procedure benefits from scale: achieving improved performance as the agent is alloted greater amounts of test-time computation. Our code and models are publicly released at jykoh.com/search-agents.
@@ 2 Background
Figure 1: Our proposed search algorithm. At each iteration, we pick the next state $s_{p}$ to expand from frontier $\mathcal{F}$ and compute a score $v$ for it using the value function. Then, we add the possible next states that the agent can get to from $s_{p}$ to the frontier, and repeat the search procedure. Faded nodes indicate explored and pruned states. Blue dashed arrows indicate backtracking.
@@ 2.1 Realistic Simulated Web Environments
Towards the goal of developing autonomous web agents powered by large language models, several prior works focused on building evaluation benchmarks for measuring the progress of models on web tasks. Mind2Web (Deng et al., 2023) is an evaluation benchmark that measures the ability of frontier models in predicting actions taken on static Internet pages. VisualWebBench (Liu et al., 2024b) introduced a multimodal benchmark for assessing the ability of models to understand web content. Others have looked towards simulators (as opposed to static HTML content): MiniWoB (Shi et al., 2017; Liu et al., 2018) was one of the first interactive simulators for web tasks, but consisted of simplified environments that do not directly translate into real world performance. WebShop (Yao et al., 2022a) simulates a simplified e-commerce site with real world data. WebLINX (Lù et al., 2024) proposes a benchmark for tackling conversational web navigation, which involves communication between the agent and a human instructor. MMInA (Zhang et al., 2024b) and OSWorld (Xie et al., 2024a) propose benchmarks to measure the ability of agents to accomplish tasks by navigating across multiple computer applications. WorkArena (Drouin et al., 2024) is a simulated environment for tasks on the ServiceNow platform. WebArena (WA) (Zhou et al., 2024b) is a benchmark of 812 tasks across 5 realistic self-hosted re-implementations of popular websites (Shopping, Reddit, CMS, GitLab, Maps), each populated with real world data. VisualWebArena (VWA) (Koh et al., 2024) is a multimodal extension to WebArena, consisting of 910 new tasks across realistic re-implementations of 3 popular real world sites (Classifieds, Reddit, Shopping). To solve tasks in VWA, agents must leverage visual grounding and understand image inputs, providing a realistic and challenging test for multimodal agents.
@@
Table 1: Possible actions $A$ in the (Visual)WebArena environments. Reproduced with permission from Koh et al. (2024).
@@ 2.2 Language-Guided Autonomous Agents
Autonomous web agents, powered by frontier (multimodal) language models (Google, 2023; OpenAI, 2024; Anthropic, 2024), are the SOTA approaches for many of the above benchmarks. Kim et al. (2024) showed that large language models can be prompted to execute computer tasks on MiniWoB++ (Liu et al., 2018), requiring far fewer demonstrations than reinforcement learning methods. AutoWebGLM (Lai et al., 2024) collects web browsing data for curriculum training and develops a web navigation agent based off a 6B parameter language model that outperforms GPT-4 on WebArena. Patel et al. (2024) showed that a language model agent can improve its performance through finetuning on its own synthetically generated data. Pan et al. (2024) show that introducing an automatic evaluator to provide guidance on task failure or success can improve the performance of a baseline Reflexion (Shinn et al., 2024) agent. Fu et al. (2024) extracts domain knowledge from offline data and provides this to the language agent during inference, to enable it to leverage helpful domain knowledge. SteP (Sodhi et al., 2024) and AWM (Wang et al., 2024) propose methods to enable agents to dynamically compose policies to solve web tasks.
In the multimodal setting, WebGUM (Furuta et al., 2024) finetuned a 3B parameter multimodal language model on a large corpus of demonstrations, achieving strong performance on MiniWoB and WebShop. Koh et al. (2024) showed that prompting multimodal language models with a Set-of-Marks (Yang et al., 2023a) representation enables the model to navigate complex webpages more effectively than text-only agents. SeeAct (Zheng et al., 2024) demonstrated that frontier multimodal models such as GPT-4V (Yang et al., 2023b) and Gemini (Google, 2023) can be grounded and prompted to solve web tasks. ICAL (Sarch et al., 2024) builds a memory of multimodal insights from demonstrations and human feedback, improving performance on VisualWebArena.
Our procedure is an inference-time approach that is compatible with many of these past approaches that focus on developing better base agents.
@@ 2.3 Search and Planning
Our method also draws inspiration from a rich history of search and planning algorithms in computer science. Search algorithms such as breadth first search, depth first search, and A* search (Hart et al., 1968) have long been used in artificial intelligence systems. Newell et al. (1959) and Laird (2019) cast goal-oriented behavior as search through a space of possible states. Dean et al. (1993) and Tash & Russell (1994) proposed planning algorithms over a limited search horizon, and employed an expansion strategy to improve plans based off heuristics about the value of information. Tash & Russell (1994) showed that this allowed agents to provide appropriate responses to time pressure and randomness in the world. Deep Blue (Campbell et al., 2002), the chess engine which defeated world champion Kasparov in chess in 1997, was based on massive parallel tree search. Pluribus (Brown & Sandholm, 2019) leverages search to find better multiplayer poker strategies for dynamic situations.
In deep learning, search algorithms with neural network components have been instrumental in achieving superhuman performance on many games: Monte-Carlo Tree Search (MCTS) (Browne et al., 2012) was used to provide lookahead search, and was pivotal in the AlphaGo (Silver et al., 2016) and AlphaGo Zero Silver et al. (2017) systems that achieved superhuman performance in the game of Go. Gray et al. (2020) performs one-step lookahead search to achieve SOTA on no-press Diplomacy. More recently, several papers (Yao et al., 2024; Besta et al., 2024) showed the potential of applying search to large language models to introduce exploration over multiple reasoning paths, enhancing performance on text based tasks that require non-trivial planning. Others have applied MCTS (Xie et al., 2024b; Chen et al., 2024a; Zhang et al., 2024a; Zhou et al., 2024a; Hao et al., 2023) to improve the performance of LMs on math benchmarks (Cobbe et al., 2021) or simplified environments (Yao et al., 2022a; Valmeekam et al., 2023).
In contrast to prior work, our setting is grounded in realistic web environments, and we search over the actual environment space (i.e., the web). This means that the search mechanics need to incorporate not just the text outputs of the agent, but also external environmental feedback from a highly complex environment.
@@ 3 Method
In this section, we describe the search procedure (Fig. 1) in detail. Successfully solving a task in a web environment such as (V)WA can be interpreted as navigating to some goal state $s_{*}$ which gives a positive reward $R(s_{*})=1$. The agent starts at state $s_{0}$ (e.g., the homepage). Given a natural language task instruction $I$, the agent’s goal is to navigate to the goal state by executing a set of actions $(a_{0},\ldots,a_{t})\in\mathcal{A}$. Each action produces a new state $s_{t+1}\in\mathcal{S}$ and observation $o_{t+1}\in\Omega$ from the environment. The transition $s_{t}\rightarrow s_{t+1}$ is governed by a deterministic transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$.
Most approaches treat this as a partially observable Markov decision process, and only condition on the current observation $o_{t}$ when predicting the next action $a_{t}$ to take. This has significant limitations: the error of the agent compounds with each step, and if an erroneous action is taken at time $t$, it cannot be easily rectified if this leads to a bad state in the future. Our approach aims to alleviate this by explicitly conducting search and backtracking to identify better trajectories. There are several components involved which we describe in the following sections: the baseline agent model (Sec. 3.1), the value function (Sec. 3.2), and the search algorithm (Sec. 3.3)
@@ 3.2 Value Function
We implement a best-first search heuristic using a value function $f_{v}$ which estimates the expected reward $\mathbb{E}[R(s_{t})]$ of the current state $s_{t}$, where the ground truth goal state would provide perfect reward of 1. As the state $s_{t}$ of the simulator is not always accessible to the agent ($s_{t}$ may include private information such as database entries of the site), the value function computes the value $v_{t}$ using the current and previous observations, as well as the natural language task instruction $I$:
@@
In our experiments, the value function is implemented by prompting a multimodal language model with the natural language instruction and observations as screenshots (Sec. 4.1).
@@ 3.3 Search Algorithm
Our proposed search algorithm is a best-first search method loosely inspired by A* search (Hart et al., 1968), a classic graph traversal algorithm used widely in computer science. We use a language model agent to propose candidate branches of the search tree. The search has hyperparameters depth $d$, branching factor $b$, and search budget $c$ which determine the maximum size of the search tree, [^1] In Sec. 5.1 we show that increasing the size of the search tree improves results at the expense of using increased compute for exploration. and termination threshold $\theta$. The search procedure is summarized in Fig. 1. We describe it in detail in the following paragraphs and provide the formal algorithm in Appendix A.4.
@@
Then, we increment the search counter $s$, and if $v_{p}$ is higher than the current best score $\hat{v}_{t}$, we update it and our best state accordingly:
@@
This concludes one iteration of search. If both of the termination conditions have not been reached, we backtrack and repeat this process for the next best state from the updated frontier $\mathcal{F}$.
@@ 4 Experiments
We run experiments on the full set of 910 VisualWebArena (VWA) and 812 WebArena (WA) tasks. These tasks are distributed across a set of diverse and realistic web environments: a Classifieds, Reddit, and Shopping environment for VWA, and a Shopping, CMS, Reddit, GitLab and Maps environment for WA.
@@ 4.1 Implementation Details
Our search algorithm is compatible with most off-the-shelf language model agents. In this work, we test it with simpler, more general, prompt-based agents, and leave incorporation of our method with more performant methods that incorporate domain-specific techniques (Fu et al., 2024; Sodhi et al., 2024) for future work. We run several prompt-based agent baselines with different input formats (full prompts provided in the appendix):
  • Multimodal SoM: For multimodal models that accept multiple image-text inputs, such as GPT-4o (OpenAI, 2024) (gpt-4o-2024-05-13), we run the multimodal agent from Koh et al. (2024) with the same prompt. We similarly apply a preprocessing step to assign a Set-of-Marks (SoM) (Yang et al., 2023a) representation to the webpage. This highlights every interactable element on the webpage with a bounding box and a unique ID. The input to the agent is a screenshot of the SoM-annotated webpage, and a text description of the elements on the page with their corresponding SoM IDs.
  • Caption-augmented: For base models that are not multimodal (e.g., Llama-3-70B-Instruct (Dubey et al., 2024)), we run the caption-augmented agent with the same prompt from Koh et al. (2024). We generate captions for each image on the webpage using an off-the-shelf captioning model (in our case, BLIP-2; Li et al. 2023). The accessibility tree [^3] https://developer.mozilla.org/en-US/docs/Glossary/Accessibility_tree representation of the webpage observation is extracted and provided to the language model as its input observation at each step.
  • Text-only: On WebArena (which does not require visual grounding), we run text-only agents using the prompt from Zhou et al. (2024b), for both GPT-4o and Llama-3-70B-Instruct. Similar to the caption-augmented baseline, this model uses an accessibility tree representation of the current webpage as its input observation (but this text-only baseline does not include image captions).
@@
We run these agents with and without search. Our search parameters are set to $d=5,b=5,c=20$, and we stop execution after a maximum of 5 actions. We enforce these constraints due to compute and budget limitations, though we expect that increasing these parameters is likely to further improve results (see Sec. 5.1 for results on scaling search parameters). We note that the fairly strict limitations on maximum actions imply that there are certain tasks that are intractable (e.g., VWA tasks with “hard” action difficulty usually require humans to execute 10 or more actions to complete). Despite this, our results show that GPT-4o with search capped at 5 max actions still substantially outperforms the GPT-4o baseline (without search) with 30 max actions.
@@
We sample actions using nucleus sampling (Holtzman et al., 2020) with a temperature of 1.0 and top-$p$ of 0.95 for all experiments. At each step of execution, we generate 20 outputs from the model by prompting it with CoT reasoning (Wei et al., 2022), and aggregate the count of the action candidates. We use the top-$b$ actions with the highest counts for branching.
@@
As detailed in Sec. 3.2, we require a value function which scores the likelihood that the current state $s_{t}$ is a goal state. We implement the value function by prompting a multimodal language model with the task instruction $I$, screenshots of the agent’s trajectory, previous actions the agent took, and the current page URL. The full prompt is provided in Appendix A.3.2. The multimodal LM is instructed to output whether the current state is a success, a failure, and if it’s a failure, whether it is on a trajectory towards success. These outputs are assigned values of 1, 0, and 0.5 respectively (and 0 for invalid output). In order to get more finegrained and reliable scores, we leverage ideas from self-consistency prompting (Wang et al., 2023), and sample multiple reasoning paths by prompting the multimodal LM with CoT (Wei et al., 2022). We sample 20 different paths from the GPT-4o model using ancestral sampling (temperature of 1.0 and top-$p$ of 1.0). The final value assigned to state $s_{t}$, used in the best-first search heuristic, is computed by averaging the values from each of the 20 reasoning paths. In our implementation, calling the value function is significantly cheaper than predicting the next action, as action prediction consumes more input tokens for few-shot examples and the representation of the page. We estimate the API cost of the GPT-4o SoM agent for action prediction to be approximately 2×2$\times$2 × that of computing the value.

### v3

@@ Abstract
Autonomous agents powered by language models (LMs) have demonstrated promise in their ability to perform decision-making tasks such as web automation. However, a key limitation remains: LMs, primarily optimized for natural language understanding and generation, struggle with multi-step reasoning, planning, and using kenvironmental feedback when attempting to solve realistic computer tasks. Towards addressing this, we propose an inference-time search algorithm for LM agents to explicitly perform exploration and multi-step planning in interactive web environments. Our approach is a form of best-first tree search that operates within the actual environment space, and is complementary with most existing state-of-the-art agents. It is the first tree search algorithm for LM agents that shows effectiveness on realistic web tasks. On the challenging VisualWebArena benchmark, applying our search algorithm on top of a GPT-4o agent yields a 39.7% relative increase in success rate compared to the same baseline without search, setting a state-of-the-art success rate of 26.4%. On WebArena, search also yields a 28.0% relative improvement over a baseline agent, setting a competitive success rate of 19.2%. Our experiments showcase the effectiveness of search for web agents, and we demonstrate that performance scales with increased test-time compute.
@@ 1 Introduction
Building agents that can perceive, plan, and act autonomously has been a long standing goal of artificial intelligence research (Russell & Norvig, 1995; Franklin & Graesser, 1996). In recent years, the advent of large language models (LMs) with strong general capabilities has paved the way towards building language-guided agents that can automate computer tasks. However, the best LM agents today are still far worse than humans. On the realistic web benchmarks WebArena (Zhou et al., 2024b) and VisualWebArena (Koh et al., 2024), humans succeed on 78% and 89% of tasks respectively, but agents — even those powered by the latest frontier models — are far worse, typically achieving success rates below 20%. One significant bottleneck in existing agents arises from their inability to leverage test-time computation for exploration and multi-step planning. Search and planning is especially important in open ended web environments, as the potential action space (i.e., all possible actions one can take on a webpage) is much larger than in most video games or text-based simulators. There are often multiple plausible actions that must be sequenced to reach a goal, and being able to efficiently explore and prune trajectories is essential. In artificial intelligence systems, one effective strategy for leveraging test-compute to improve results is search: iteratively constructing, exploring, and pruning a graph of intermediate states and possible solutions (Newell et al., 1959; Silver et al., 2016; Laird, 2019). The effectiveness of search algorithms has been shown time and time again, enabling models to achieve or surpass human-level performance on a variety of games, including Go (Silver et al., 2016; 2017), poker (Brown & Sandholm, 2018; 2019), and Diplomacy (Gray et al., 2020).
Figure 1: Our proposed search algorithm. At each iteration, we pick the next state $s_{p}$ to expand from frontier $\mathcal{F}$ and compute a score $v$ for it using the value function. Then, we add the possible states that the agent can get to from $s_{p}$ to the frontier and repeat the search procedure. Faded nodes indicate explored and pruned states. Blue dashed arrows indicate backtracking.
@@
Our experiments show that this search procedure is complementary with existing LM agents, and enables these models to perform better on harder and longer horizon tasks. On VisualWebArena (Koh et al., 2024), search improves the performance of a baseline GPT-4o (OpenAI, 2024) agent by 39.7% relative to the baseline without search, setting a new state-of-the-art (SOTA) success rate of 26.4%. On WebArena (Zhou et al., 2024b), search is also highly effective, contributing a 28.0% relative improvement (yielding a competitive success rate of 19.2%). We also demonstrate that our search procedure benefits from scale: achieving improved performance as the agent is allotted greater amounts of test-time computation. Our code and models are publicly released at jykoh.com/search-agents.
@@ 2.1 Realistic Simulated Web Environments
Towards the goal of developing autonomous web agents powered by large language models, several prior works focused on building evaluation benchmarks for measuring the progress of models on web tasks. Mind2Web (Deng et al., 2023) is an evaluation benchmark that measures the ability of frontier models in predicting actions taken on static Internet pages. VisualWebBench (Liu et al., 2024b) introduced a multimodal benchmark for assessing the ability of models to understand web content. Others have looked towards simulators (as opposed to static HTML content): MiniWoB (Shi et al., 2017; Liu et al., 2018) was one of the first interactive simulators for web tasks, but consisted of simplified environments that do not directly translate into real world performance. WebShop (Yao et al., 2022a) simulates a simplified e-commerce site with real world data. WebLINX (Lù et al., 2024) proposes a benchmark for tackling conversational web navigation, which involves communication between the agent and a human instructor. MMInA (Zhang et al., 2024c) and OSWorld (Xie et al., 2024a) propose benchmarks to measure the ability of agents to accomplish tasks by navigating across multiple computer applications. WorkArena (Drouin et al., 2024) is a simulated environment for tasks on the ServiceNow platform. Outside of the web, environments such AndroidWorld (Rawles et al., 2024) is a dynamic environment for measuring the performance of agents on mobile applications. Several other work build comprehensive, highly realistic, fully-featured web simulators. WebArena (WA) (Zhou et al., 2024b) is a benchmark of 812 tasks across 5 realistic self-hosted re-implementations of popular websites (Shopping, Reddit, CMS, GitLab, Maps), each populated with real world data. VisualWebArena (VWA) (Koh et al., 2024) is a multimodal extension to WebArena, consisting of 910 new tasks across realistic re-implementations of 3 popular real world sites (Classifieds, Reddit, Shopping). To solve tasks in VWA, agents must leverage visual grounding and understand image inputs — a realistic and challenging test for multimodal agents.
@@
Table 1: Possible actions AA in the (Visual)WebArena environments.
@@ 2.2 Language-Guided Autonomous Agents
Autonomous web agents, powered by frontier (multimodal) language models (Google, 2023; OpenAI, 2024; Anthropic, 2024), are the SOTA approaches for many of the above benchmarks. Kim et al. (2024) showed that large language models can be prompted to execute computer tasks on MiniWoB++ (Liu et al., 2018), requiring far fewer demonstrations than reinforcement learning methods. AutoWebGLM (Lai et al., 2024) collects web browsing data for curriculum training and develops a web navigation agent based off a 6B parameter language model that outperforms GPT-4 on WebArena. Patel et al. (2024) showed that a language model agent can improve its performance through finetuning on its own synthetically generated data. Pan et al. (2024) show that introducing an automatic evaluator to provide guidance on task failure or success can improve the performance of a baseline Reflexion (Shinn et al., 2024) agent. Fu et al. (2024) extracts domain knowledge from offline data and provides this to the language agent during inference, to enable it to leverage helpful domain knowledge. SteP (Sodhi et al., 2024) and AWM (Wang et al., 2024b) propose methods to enable agents to dynamically compose policies to solve web tasks.
In the multimodal setting, WebGUM (Furuta et al., 2024) finetuned a 3B parameter multimodal language model on a large corpus of demonstrations, achieving strong performance on MiniWoB and WebShop. Koh et al. (2024) showed that prompting multimodal language models with a Set-of-Marks (Yang et al., 2023) representation enables the model to navigate complex webpages more effectively than text-only agents. SeeAct (Zheng et al., 2024) demonstrated that frontier multimodal models can be grounded and prompted to solve web tasks. ICAL (Sarch et al., 2024) builds a memory of multimodal insights from demonstrations and human feedback. Our procedure is an inference-time approach that is compatible with many of these past approaches that focus on developing better base agents.
@@ 2.3 Search and Planning
Our method also draws inspiration from a rich history of search and planning algorithms in computer science. Search algorithms such as breadth-first search, depth-first search, and A* search (Hart et al., 1968) have long been used in artificial intelligence systems. Newell et al. (1959) and Laird (2019) cast goal-oriented behavior as search through a space of possible states. Dean et al. (1993) and Tash & Russell (1994) proposed planning algorithms over a limited search horizon, and employed an expansion strategy to improve plans based off heuristics. Tash & Russell (1994) showed that this allowed agents to provide appropriate responses to time pressure and randomness in the world. Deep Blue (Campbell et al., 2002), the chess engine which defeated world champion Kasparov in chess in 1997, was based on massive parallel tree search. Pluribus (Brown & Sandholm, 2019) leverages search to find better multiplayer poker strategies for dynamic situations.
In deep learning, search algorithms with neural network components have been instrumental in achieving superhuman performance in many games. Monte-Carlo Tree Search (MCTS) (Browne et al., 2012) was used to provide lookahead search in the AlphaGo (Silver et al., 2016; 2017) systems that achieved superhuman performance in the game of Go. Gray et al. (2020) performs one-step lookahead search to achieve SOTA on no-press Diplomacy. More recently, several papers (Yao et al., 2024; Besta et al., 2024) showed the potential of applying search to large language models to introduce exploration, enhancing performance on text based tasks that require non-trivial planning. Others have applied MCTS (Hao et al., 2023; Zhou et al., 2024a; Xie et al., 2024b; Chen et al., 2024a; Zhang et al., 2024b; Wang et al., 2024a; Zhang et al., 2024a) to improve the performance of LMs on math and science benchmarks (Cobbe et al., 2021; Wang et al., 2023a) or simplified environments (Yao et al., 2022a; Valmeekam et al., 2023; Zhou et al., 2024a).
In contrast to prior work, we search over the actual environment space of realistic, complex websites. This means that search mechanics need to incorporate not just the text outputs of the agent, but also external environmental feedback.
@@ 3 Method
In this section, we describe the search procedure (Fig. 1) in detail. Successfully solving a task in a web environment such as (V)WA can be interpreted as navigating to a goal state $s_{*}$ which gives a positive reward $R(s_{*})=1$. The agent starts at state $s_{0}$ (e.g., the homepage). Given a natural language instruction $I$, the agent’s goal is to navigate to $s_{*}$ by executing actions $(a_{0},\ldots,a_{t})\in\mathcal{A}$. Each action produces a new state $s_{t+1}\in\mathcal{S}$ and observation $o_{t+1}\in\Omega$ from the environment. The transition $s_{t}\rightarrow s_{t+1}$ is governed by a deterministic transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$.
Most approaches treat this as a partially observable Markov decision process, and only condition on the current observation $o_{t}$ when predicting the next action $a_{t}$ to take. This has significant limitations: the error of the agent compounds with each step, and if an erroneous action is taken at time $t$, it cannot be easily rectified if this leads to a bad state. Our approach aims to alleviate this by explicitly conducting search and backtracking to identify better trajectories.
@@ 3.2 Value Function
We implement a best-first search heuristic using a value function $f_{v}$ which estimates the expected reward $\mathbb{E}[R(s_{t})]$ of the current state $s_{t}$, where the ground truth goal state would provide perfect reward of 1. As the state $s_{t}$ of the simulator is not always accessible to the agent ($s_{t}$ may include private information such as site database entries), the value function computes the value $v_{t}$ using the current and previous observations, as well as the natural language task instruction $I$:
@@
In our experiments (Sec. 4.1), the value function is implemented by prompting a multimodal language model with the task instruction and observation screenshots.
@@ 3.3 Search Algorithm
Our proposed search algorithm is a best-first search method loosely inspired by A* search (Hart et al., 1968), a classic graph traversal algorithm used widely in computer science. We use a language model agent to propose candidate branches of the search tree. The search has hyperparameters depth $d$, branching factor $b$, and search budget $c$ which determine the maximum size of the search tree, [^1] In Sec. 5.1 we show that increasing the size of the search tree improves results by leveraging increased compute. and termination threshold $\theta$. The search procedure is summarized in Fig. 1. We describe it in detail in the following paragraphs and provide the formal algorithm in Appendix A.4.
@@
Then, we increment $s$, and if $v_{p}$ is higher than the current best score $\hat{v}_{t}$, we update it and our best state accordingly:
@@
This concludes one iteration of search. If both termination conditions have not been reached, we backtrack and repeat this for the next best state from the updated frontier $\mathcal{F}$.
@@ 4 Experiments
We run experiments on the full set of 910 VisualWebArena (VWA) and 812 WebArena (WA) tasks. The tasks are distributed across a set of diverse and realistic websites.
@@ 4.1 Implementation Details
Our search algorithm is compatible with most off-the-shelf language model agents. In this work, we test it with simpler, more general, prompt-based agents, and leave incorporation of our method with more performant methods that incorporate domain-specific techniques (Fu et al., 2024; Sodhi et al., 2024) for future work. We run several prompt-based agent baselines:
  • Multimodal SoM: For multimodal models that accept multiple image-text inputs, such as GPT-4o (OpenAI, 2024) (gpt-4o-2024-05-13), we run the multimodal agent from Koh et al. (2024) with the same prompt. We similarly apply a preprocessing step to assign a Set-of-Marks (SoM) (Yang et al., 2023) representation to the webpage. This highlights every interactable element on the webpage with a bounding box and a unique ID. The input to the agent is a screenshot of the SoM-annotated webpage, and a text description of the elements on the page with their assigned IDs.
  • Caption-augmented: For base models that are not multimodal (e.g., Llama-3-70B-Instruct (Dubey et al., 2024)), we run the caption-augmented agent with the same prompt from Koh et al. (2024). We generate captions for each image on the webpage using an off-the-shelf captioning model (in our case, BLIP-2; Li et al. 2023). The accessibility tree [^3] https://developer.mozilla.org/en-US/docs/Glossary/Accessibility_tree representation of the webpage is used as the input observation.
  • Text-only: On WebArena (which does not require visual grounding), we run text-only agents using the prompt from Zhou et al. (2024b), for both GPT-4o and Llama-3-70B-Instruct. This model uses an accessibility tree (w/o captions) of the current page as input.
@@
Our search parameters are set to $d=5,b=5,c=20$, and we stop execution after a maximum of 5 actions. We enforce these constraints due to compute and budget limitations, though we expect that increasing these parameters is likely to further improve results (see Sec. 5.1 for results on scaling search parameters). We note that the fairly strict limitations on maximum actions imply that there are certain tasks that are intractable (e.g., VWA tasks with “hard” action difficulty usually require humans to execute 10 or more actions to complete). Despite this, our results show that GPT-4o with search capped at 5 max actions still substantially outperforms the GPT-4o baseline (without search) with 30 max actions.
@@
We sample actions using nucleus sampling (Holtzman et al., 2020) with a temperature of 1.0 and top-$p$ of 0.95 for all experiments. At each step of execution, we generate 20 outputs from the model by prompting it with CoT reasoning (Wei et al., 2022). We aggregate the count of the actions and use the top-$b$ actions for branching.
@@
As detailed in Sec. 3.2, we require a value function which scores the likelihood that the current state $s_{t}$ is a goal state. We implement the value function by prompting a multimodal language model with the task instruction $I$, screenshots of the agent’s trajectory, previous actions the agent took, and the current page URL. The full prompt is provided in Appendix A.3.2. The multimodal LM is instructed to output whether the current state is a success, a failure, and if it’s a failure, whether it is on a trajectory towards success. These outputs are assigned values of 1, 0, and 0.5 respectively (and 0 for invalid output). In order to get more finegrained and reliable scores, we leverage ideas from self-consistency prompting (Wang et al., 2023b), and sample multiple reasoning paths by prompting the multimodal LM with CoT (Wei et al., 2022). We sample 20 different paths from the GPT-4o model using ancestral sampling (temperature of 1.0 and top-$p$ of 1.0). The final value assigned to state $s_{t}$, used in the best-first search heuristic, is computed by averaging the values from each of the 20 reasoning paths. In our implementation, calling the value function is significantly cheaper than predicting the next action, as action prediction consumes more input tokens for few-shot examples and the representation of the page.  [^4] We estimate the API cost of the GPT-4o SoM agent for action prediction to be approximately $2\times$ that of computing the value.